OpenStudy (anonymous):

Write y-5=-(x-7) in standard form

OpenStudy (anonymous):

y-5 =-(x-7) y-5 = -x + 7 x+y -5 -7 = 0 x+ y -12 = 0

OpenStudy (anonymous):

Wait why are we adding 7

OpenStudy (anonymous):

we are multiplying - sign to x - 7